OpenStudy (anonymous):

\[2x ^{1/2}-x ^{1/4}=3\]

OpenStudy (dumbcow):

its really just a quadratic u = x^1/4 \[2u ^{2}-u-3 =0\] factor \[(2u-3)(u+1) = 0\] u = 3/2, -1 x = u^4 x=(-1)^4 = 1 x = (3/2)^4 = 81/16 check for extraneous solutions 2(1) - 1 not equal 3 solution: x = 81/16